# Durham, UK Image Inpainting

### v1

@@ Abstract
Image inpainting, or image completion, is a crucial task in computer vision that aims to restore missing or damaged regions of images with semantically coherent content. This technique requires a precise balance of local texture replication and global contextual understanding to ensure the restored image integrates seamlessly with its surroundings. Traditional methods using Convolutional Neural Networks (CNNs) are effective at capturing local patterns but often struggle with broader contextual relationships due to the limited receptive fields. Recent advancements have incorporated transformers, leveraging their ability to understand global interactions. However, these methods face computational inefficiencies and struggle to maintain fine-grained details. To overcome these challenges, we introduce MxT composed of the proposed Hybrid Module (HM), which combines Mamba with the transformer in a synergistic manner. Mamba is adept at efficiently processing long sequences with linear computational costs, making it an ideal complement to the transformer for handling long-scale data interactions. Our HM facilitates dual-level interaction learning at both pixel and patch levels, greatly enhancing the model to reconstruct images with high quality and contextual accuracy. We evaluate MxT on the widely-used CelebA-HQ and Places2-standard datasets, where it consistently outperformed existing state-of-the-art methods.
@@ 1 Introduction
Convolutional Neural Networks (CNNs) have been employed for image inpainting, capitalizaing on their ability to capture local patterns and textures. However, CNNs-based methods are inherently limited by the slow grown receptive field, which limit the ability to grasp broader image context [Li et al.(2022a)Li, Lin, Zhou, Qi, Wang, and Jia, Zamir et al.(2022)Zamir, Arora, Khan, Hayat, Khan, and Yang]. To solve this issue, recent advancements [Li et al.(2022a)Li, Lin, Zhou, Qi, Wang, and Jia, Chen et al.(2024)Chen, Atapour-Abarghouei, and Shum] have seen the integration of transformer or self-attention into image inpainting, leveraging their capability to capture global correlations across entire images. However, transformer-based methods are often constrained by quadratic computational complexity, prompting most methods to process images in smaller patches to reduce the spatial dimension [Wan et al.(2021)Wan, Zhang, Chen, and Liao, Zheng et al.(2022)Zheng, Cham, Cai, and Phung], to learn the interaction in patch-level. This patch-based approach hinder the learning of fine-grained details, often resulting in artifacts in the generated images.
@@
We observe that, Mamba and transformer exhibit complementary strengths: Mamba is good at learning long-range pixel-wise dependency, which is computationally expensive for the transformer. Conversely, transformer is good at in capturing global interactions between localized patches, such spatial awareness is an area where Mamba lacks due to it is designed for sequence modelling.
In this paper, we introduce MxT, consisted by proposed Hybrid Modules that synergistically combines the strengths of both transformer and Mamba. This novel approach allows for a dual-level interaction learning from the patch level and pixel level. Our comparative experiments demonstrate that our MxT outperforms existing state-of-the-art methods on two wildly used datasets, CelebA-HQ and Places2. We summarize our contributions as: 1) We propose MxT to introduce Mamba conmbined with transformer for image inpainting. 2) We design a novel Hybrid Module to capture the feature interaction in both pixel-level and patch-level. 3) Our MxT overall suppress the state-of-the-art methods on both CelebA-HQ and Places2 datasets. 4) MxT is able to adapt to high-resolution images with only training on low-resolution data.
@@ 2.2 SSMs in Computer Vision
Recently, State Space Models (SSMs) have demonstrated promising advantages of long sequence modeling and linear-time complexity in Natural Language Processing (NLP) [Gu et al.(2021)Gu, Goel, and Re]. This work specifically tackles the problem of vanishing gradients in SSMs when solving the exponential function by the linear first-order Ordinary Differential Equations [Gu et al.(2020)Gu, Dao, Ermon, Rudra, and Ré]. Building opon the rigorous theoretical proofs of the HiPPO framework that enables SSMs to capture long-range dependencies, Gu et al. [Gu and Dao(2023)] further introduce a data-dependent selective structure SSM (i.e., Mamba) to significantly improve the computational efficiencies in conventional SSMs. Inspired by pioneering SSMs, vision-specific adaptations of the Mamba architecture, such as Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and V-Mamba [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], propose visual SSMs designs for computer vision tasks including image classification and object detection [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. However, their performance are still behind of the state-of-the-art transformer-based models like SpectFormer [Patro et al.(2023)Patro, Namboodiri, and Agneeswaran], SVT [Patro and Agneeswaran(2024)], and WaveViT [Yao et al.(2022)Yao, Pan, Li, Ngo, and Mei]. U-Mamba [Ma et al.(2024)Ma, Li, and Wang] effectively extends the capabilities of Mamba for biomedical image segmentation by proposing a hybrid CNN-SSM block. However, these studies ineffectively leverage the capabilities of Mamba in image long-range pixel-level dependency learning and overlook the critical spatial awareness during model designs.
@@ 3.1 State Space Models and Mamba
where $\textbf{{A}}\in\mathbb{R}^{N\times N}$, $\textbf{{B}}\in\mathbb{R}^{N\times 1}$, $\textbf{{C}}\in\mathbb{R}^{1\times N}$, and $\Delta$ is a time-scale parameter. Mamba [Gu and Dao(2023)], one of the most recent selective SSMs, introduces a gated selective mechanism to propagate or eliminate selected information based on the current state, significantly improves the content-reasoning performance. Specifically, Mamba change the model from time-invariant to time-varying via converting parameters $\Delta$, B, C into input-dependent functions.
Figure 1: (a) The architecture overview of the proposed MxT. (b) The Hybrid Block composed of $n$ proposed Hybrid Modules. (c) The proposed Hybrid Module, consisted by a Mamba Block, a Spatial Reduced Self-Attention and a Context Broadcasting Feed-forward Network. (d) The Spatial Reduced Self-Attention provides spatial awareness. (e) The Mamba Block captures pixel-level interaction. (f) The Context Broadcasting Feed-forward Network transfers the features.
@@ 4 Method
The overall pipeline of the proposed MxT is illustrated in Fig. 1, which is a U-Net shape architecture formed with 7 Hybrid Blocks. Formally, the masked image ${I}_{masked}\in\mathbb{R}^{H\times W\times 3}$ concatenated with a mask ${M}\in\mathbb{R}^{H\times W\times 1}$ as the input ${I}_{in}$. We first use a overlapped convolution to embed ${I}_{in}$, then feed into the following 7 HBs with 3 times downsampling and 3 times upsampling. At the end, one convolution layer project the final output ${I}_{out}$. Each Hybrid Block is consisted with $n$ Hybrid Modules (HM), as shown in Fig. 1 (b), where $n$ is the number of HMs. Each HM has a Transformer block, a Mamba block and a Context Broadcasting Feedforward Network (CBFN), which will be detailed in the section 4.1.
@@ 4.1 Hybrid Module
Each of the seven Hybrid Modules involving a pair of SRSA (Spatial Reduced Self-Attention) and Mamba modules for capturing long-range dependency, followed by a Context Broadcasting Feedforward Network (CBFN) to enhance the local context and control data flow consistency.
Spatial Reduced Self-Attention. We introduce the Spatial Reduced Self-Attention (SRSA) module, designed to leverage the capability of transformer for capturing global correlation while enriching local context detail. Specifically, given a input feature $F$, we first apply layer normalization followed by a $1\times 1$ convolution and a $3\times 3$ depth-wise convolution to extract the local features:
@@
Mamba with Positional Embedding. Mamba showcase the strong capacity in handling long sequence data with linear computational complexity, making it highly effective for modeling interactions between adjacent pixels. In this work, we propose leveraging Mamba module to modeling the flattened feature, thereby capturing long-range dependency at the pixel level, which is expensive to capture by self-attention. To adapt Mamba more aptly for vision task and enhance its ability to maintain positional awareness, we incorporate positional embedding into the module.
@@
Subsenquently, we introduce cosine positional embedding [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] to the transformed feature, enhancing the capacity to maintain positional awareness,:
@@
The gate branch involves a linear layer and a SiLU activation function [Hendrycks and Gimpel(2016)]. After the gate branch re-weight the body branch, the output will be reshape to the shape of $(B,C,H,W)$:
@@
Context Broadcasting Feed-forward Network. We propose Context Broadcasting Feedforward Network (CBFN) by improving the Gated-Dconv Feed-Forward Network (GDFN) [Zamir et al.(2022)Zamir, Arora, Khan, Hayat, Khan, and Yang]. The GDFN, recognized for its efficacy in enhancing local context through a gated mechanism with depth-wise convolution. To build upon this, our CBFN integrates a global processing stage post-GDFN. Specifically, we implement global average pooling followed by channel-wise averaging to obtain the overall mean value of the input feature $F$, denoted as $\mu=GlobalAvgPool(F)$, where $F$ is the output from GDFN. This $\mu$ is then broadcast to the dimensions of $F$ and added to it. The output of CBFN is represented as $F^{\prime}$:
@@ 5 Experiment Results
Datasets. We evaluate our MxT on two diverse datasets, CelebA-HQ [Karras et al.(2017)Karras, Aila, Laine, and Lehtinen] and Places2-standard [Zhou et al.(2017)Zhou, Lapedriza, Khosla, Oliva, and Torralba], to ensure a comprehensive comparison. CelebA-HQ is a dataset consisting of high-quality human face images. For CelebA-HQ, we train our model on the first 28000 images and reserve the remaining 2000 for testing. Places2 comprises a wide range of natural and indoor scene images. For Places2, we employ the standard training set, which includes 1.8 million images, and test on its validation set of 30000 images. We follow [Guo et al.(2021)Guo, Yang, and Huang, Li et al.(2022b)Li, Guo, Lin, Li, Feng, and Wang, Chen et al.(2024)Chen, Atapour-Abarghouei, and Shum] to conduct all experiments with the widely used irregular mask [Liu et al.(2018a)Liu, Reda, Shih, Wang, Tao, and Catanzaro] in three mask ratios.
@@ 5.1 Comparison with State of the Art
Quantitative Comparison. For a fair comparison, we employ the officially released models and test them with the same test sets and masks. As shown in table 1, our MxT output perform in all metrics cross different mask ratios. Especially on CelebA-HQ, at the increasing mask ratios, MxT improves PSNR by 2.0%, 2.3% and 1.8% respectively, and decreases LPIPS by 12.3%, 11.6% and 10.5% respectively. Moreover, for the Places2, our model demonstrated comparable performance to SOTAs such as MISF and LAMA. While our training utilized the Places2-Standard dataset with 1.8 million images, MISF and LAMA were trained on the Places2-Challenge dataset, which contains 8 million images. Despite using only 22.5% of the images employed by these benchmarks, our model achieved comparable results, showcasing its robustness and efficiency.
Qualitative Comparison. The qualitative comparisons are presented in Fig. 2. For human face samples, MxT maintains consistency from the visible regions to the missing regions, such as effectively reconstructing elements like a missing hat. Additionally, MxT renders features like eyes with improved fine-grained details, showcasing its strong capability in learning complex representations. In the Places2 dataset, MxT effectively captures the spatial layouts in indoor environments and excels at maintaining the architectural integrity of the road surfaces. Such examples highlight our MxT has superior spatial perceptions.
Efficiency Comparison. Our MxT efficiently reduces spatial dimensions within the transformer and leverages Mamba’s inherent capabilities, both achieving linear complexity. This synergy optimizes operational simplicity while enhancing effectiveness. As shown in Tab. 3, our model comprises 180 million parameters and achieves 110 ms to inference one image, making it suitable for real-time applications.
@@
Table 2: Ablation studies of each component. MB is the Mamba Block with positonal embedding. SRSA is the Spatial Reduced Self-Attention. GDFN is the feed-forward network in [Zamir et al.(2022)Zamir, Arora, Khan, Hayat, Khan, and Yang]. CBFN is the Context Broadcasting Feed-forward Network. Our MxT corresponds to configuration (e).
@@ 5.2 Ablation Study
In our comprehensive ablation study conducted on CelebA-HQ, we incrementally enhance the baseline U-Net shape model, observing significant performance improvements with the integration of each component. Results are shown in the Tab. 5.1. The addition of the Mamba Block in configuration (b) and the addition of self-attention in configuration (c) both demonstrated the improvement across all evaluated metrics compared to the baseline (a). Notably, self-attention improves in SSIM, suggesting its superior capability in capture spatial interactions. Mamba showcase the superior in capture pixel-level interactions, demonstrated from the better PSNR and L1 values. The simultaneous use of Mamba and self-attention in configuration (d) lead to further improvements, indicating that these components effectively complement each other and ontribute to a robust model. Configuration (e) is our final model, where we optimize GDFN to CBFN. The overall metrics are further improved. All ablation experiments are trained for 30K iterations. In addition, we followed  [Cui et al.(2022)Cui, Tao, Bing, Ren, Gao, Cao, Huang, and Knoll] to build a light MxT with halved parameter for an efficient evaluation.
5.3 Application: High Resolution Image Inpainting
Our MxT is designed with linear computational complexity, enabling it to effectively handle high-resolution image inpainting task. We directly apply our model, pre-trained on Places2-standard dataset, to real-world high-resolution images, demonstrating its capability. The example is illustrated in Fig. 3.

### v2

# Durham, UK Mamba $\times$ Transformer for Image Inpainting

@@ Abstract
Image inpainting, or image completion, is a crucial task in computer vision that aims to restore missing or damaged regions of images with semantically coherent content. This technique requires a precise balance of local texture replication and global contextual understanding to ensure the restored image integrates seamlessly with its surroundings. Traditional methods using Convolutional Neural Networks (CNNs) are effective at capturing local patterns but often struggle with broader contextual relationships due to the limited receptive fields. Recent advancements have incorporated transformers, leveraging their ability to understand global interactions. However, these methods face computational inefficiencies and struggle to maintain fine-grained details. To overcome these challenges, we introduce M×$\times$×T composed of the proposed Hybrid Module (HM), which combines Mamba with the transformer in a synergistic manner. Mamba is adept at efficiently processing long sequences with linear computational costs, making it an ideal complement to the transformer for handling long-scale data interactions. Our HM facilitates dual-level interaction learning at both pixel and patch levels, greatly enhancing the model to reconstruct images with high quality and contextual accuracy. We evaluate M×$\times$×T on the widely-used CelebA-HQ and Places2-standard datasets, where it consistently outperformed existing state-of-the-art methods.
@@ 1 Introduction
Convolutional Neural Networks (CNNs) have been employed for image inpainting, capitalizing on their ability to capture local patterns and textures. However, CNNs-based methods are inherently limited by the slow-grown receptive field, which limits the ability to grasp broader image context [Li et al.(2022a)Li, Lin, Zhou, Qi, Wang, and Jia, Zamir et al.(2022)Zamir, Arora, Khan, Hayat, Khan, and Yang]. To solve this issue, recent advancements [Li et al.(2022a)Li, Lin, Zhou, Qi, Wang, and Jia, Chen et al.(2024)Chen, Atapour-Abarghouei, and Shum] have seen the integration of transformer or self-attention into image inpainting, leveraging their capability to capture global correlations across entire images. However, transformer-based methods are often constrained by quadratic computational complexity, prompting most methods to process images in smaller patches to reduce the spatial dimension [Wan et al.(2021)Wan, Zhang, Chen, and Liao, Zheng et al.(2022)Zheng, Cham, Cai, and Phung], to learn the interaction in patch-level. This patch-based approach hinders the learning of fine-grained details, often resulting in artifacts in the generated images.
@@
We observe that, Mamba and transformer exhibit complementary strengths: Mamba is good at learning long-range pixel-wise dependency, which is computationally expensive for the transformer. Conversely, transformer is good at capturing global interactions between localized patches, such spatial awareness is an area that Mamba lacks due to it being designed for sequence modelling.
In this paper, we introduce M×$\times$×T, consisted by proposed Hybrid Modules that synergistically combine the strengths of both transformer and Mamba. This novel approach allows for a dual-level interaction learning from the patch level and pixel level. Our comparative experiments demonstrate that our M×$\times$×T outperforms existing state-of-the-art methods on two wildly used datasets, CelebA-HQ and Places2. We summarize our contributions as: 1) We propose M×$\times$×T to introduce Mamba combined with transformer for image inpainting. 2) We design a novel Hybrid Module to capture the feature interaction in both pixel-level and patch-level. 3) Our M×$\times$×T overall suppress the state-of-the-art methods on both CelebA-HQ and Places2 datasets. 4) M×$\times$×T is able to adapt to high-resolution images with only training on low-resolution data.
@@ 2.2 SSMs in Computer Vision
Recently, State Space Models (SSMs) have demonstrated promising advantages of long sequence modeling and linear-time complexity in Natural Language Processing (NLP) [Gu et al.(2021)Gu, Goel, and Re]. This work specifically tackles the problem of vanishing gradients in SSMs when solving the exponential function by the linear first-order Ordinary Differential Equations [Gu et al.(2020)Gu, Dao, Ermon, Rudra, and Ré]. Building on the rigorous theoretical proofs of the HiPPO framework that enables SSMs to capture long-range dependencies, Gu et al. [Gu and Dao(2023)] further introduce a data-dependent selective structure SSM (i.e., Mamba) to significantly improve the computational efficiencies in conventional SSMs. Inspired by pioneering SSMs, vision-specific adaptations of the Mamba architecture, such as Vision Mamba [Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang] and V-Mamba [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu], propose visual SSMs designs for computer vision tasks including image classification and object detection [Liu et al.(2024)Liu, Tian, Zhao, Yu, Xie, Wang, Ye, and Liu, Zhu et al.(2024)Zhu, Liao, Zhang, Wang, Liu, and Wang]. However, their performance are still behind of the state-of-the-art transformer-based models like SpectFormer [Patro et al.(2023)Patro, Namboodiri, and Agneeswaran], SVT [Patro and Agneeswaran(2024)], and WaveViT [Yao et al.(2022)Yao, Pan, Li, Ngo, and Mei]. U-Mamba [Ma et al.(2024)Ma, Li, and Wang] effectively extends the capabilities of Mamba for biomedical image segmentation by proposing a hybrid CNN-SSM block. However, these studies ineffectively leverage the capabilities of Mamba in image long-range pixel-level dependency learning and overlook the critical spatial awareness during model designs.
@@ 3.1 State Space Models and Mamba
where $\textbf{{A}}\in\mathbb{R}^{N\times N}$, $\textbf{{B}}\in\mathbb{R}^{N\times 1}$, $\textbf{{C}}\in\mathbb{R}^{1\times N}$, and $\Delta$ is a time-scale parameter. Mamba [Gu and Dao(2023)], one of the most recent selective SSMs, introduces a gated selective mechanism to propagate or eliminate selected information based on the current state, significantly improves the content-reasoning performance. Specifically, Mamba changes the model from time-invariant to time-varying via converting parameters $\Delta$, B, C into input-dependent functions.
Figure 1: (a) The architecture overview of the proposed M×$\times$×T. (b) The Hybrid Block is composed of $n$ proposed Hybrid Modules. (c) The proposed Hybrid Module, consisted of a Mamba Block, a Spatial Reduced Self-Attention and a Context Broadcasting Feed-forward Network. (d) The Spatial Reduced Self-Attention provides spatial awareness. (e) The Mamba Block captures pixel-level interaction. (f) The Context Broadcasting Feed-forward Network transfers the features.
@@ 4 Method
The overall pipeline of the proposed M×$\times$×T is illustrated in Fig. 1, which is a U-Net shape architecture formed with 7 Hybrid Blocks. Formally, the masked image ${I}_{masked}\in\mathbb{R}^{H\times W\times 3}$ concatenated with a mask ${M}\in\mathbb{R}^{H\times W\times 1}$ as the input ${I}_{in}$. We first use an overlapped convolution to embed ${I}_{in}$, then feed into the following 7 HBs with 3 times downsampling and 3 times upsampling. At the end, one convolution layer projects the final output ${I}_{out}$. Each Hybrid Block is consisted of $n$ Hybrid Modules (HM), as shown in Fig. 1 (b), where $n$ is the number of HMs. Each HM has a Transformer block, a Mamba block and a Context Broadcasting Feedforward Network (CBFN), which will be detailed in the section 4.1.
@@ 4.1 Hybrid Module
Each of the seven Hybrid Modules involves a pair of SRSA (Spatial Reduced Self-Attention) and Mamba modules for capturing long-range dependency, followed by a Context Broadcasting Feedforward Network (CBFN) to enhance the local context and control data flow consistency.
Spatial Reduced Self-Attention. We introduce the Spatial Reduced Self-Attention (SRSA) module, designed to leverage the capability of the transformer for capturing global correlation while enriching local context detail. Specifically, given a input feature $F$, we first apply layer normalization followed by a $1\times 1$ convolution and a $3\times 3$ depth-wise convolution to extract the local features:
@@
Mamba with Positional Embedding. Mamba showcases a strong capacity to handle long sequence data with linear computational complexity, making it highly effective for modeling interactions between adjacent pixels. In this work, we propose leveraging the Mamba module to modeling the flattened feature, thereby capturing long-range dependency at the pixel level, which is expensive to capture by self-attention. To adapt Mamba more aptly for vision tasks and enhance its ability to maintain positional awareness, we incorporate positional embedding into the module.
@@
Subsequently, we introduce cosine positional embedding [Vaswani et al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones, Gomez, Kaiser, and Polosukhin] to the transformed feature, enhancing the capacity to maintain positional awareness:
@@
The gate branch involves a linear layer and a SiLU activation function [Hendrycks and Gimpel(2016)]. After the gate branch re-weight the body branch, the output will be reshaped to the shape of $(B,C,H,W)$:
@@
Context Broadcasting Feed-forward Network. We propose Context Broadcasting Feedforward Network (CBFN) by improving the Gated-Dconv Feed-Forward Network (GDFN) [Zamir et al.(2022)Zamir, Arora, Khan, Hayat, Khan, and Yang]. The GDFN is recognized for its efficacy in enhancing local context through a gated mechanism with depth-wise convolution. To build upon this, our CBFN integrates a global processing stage post-GDFN. Specifically, we implement global average pooling followed by channel-wise averaging to obtain the overall mean value of the input feature $F$, denoted as $\mu=GlobalAvgPool(F)$, where $F$ is the output from GDFN. This $\mu$ is then broadcast to the dimensions of $F$ and added to it. The output of CBFN is represented as $F^{\prime}$:
@@ 5 Experiment Results
Datasets. We evaluate our M×$\times$×T on two diverse datasets, CelebA-HQ [Karras et al.(2017)Karras, Aila, Laine, and Lehtinen] and Places2-standard [Zhou et al.(2017)Zhou, Lapedriza, Khosla, Oliva, and Torralba], to ensure a comprehensive comparison. CelebA-HQ is a dataset consisting of high-quality human face images. For CelebA-HQ, we train our model on the first 28000 images and reserve the remaining 2000 for testing. Places2 comprises a wide range of natural and indoor scene images. For Places2, we employ the standard training set, which includes 1.8 million images, and test on its validation set of 30000 images. We follow [Guo et al.(2021)Guo, Yang, and Huang, Li et al.(2022b)Li, Guo, Lin, Li, Feng, and Wang, Chen et al.(2024)Chen, Atapour-Abarghouei, and Shum] to conduct all experiments with the widely used irregular mask [Liu et al.(2018a)Liu, Reda, Shih, Wang, Tao, and Catanzaro] in three mask ratios.
@@ 5.1 Comparison with State of the Art
Quantitative Comparison. For a fair comparison, we employ the officially released models and test them with the same test sets and masks. As shown in table 1, our M×$\times$×T output perform in all metrics cross different mask ratios. Especially on CelebA-HQ, at the increasing mask ratios, M×$\times$×T improves PSNR by 2.0%, 2.3% and 1.8% respectively, and decreases LPIPS by 12.3%, 11.6% and 10.5% respectively. Moreover, for the Places2, our model demonstrated comparable performance to SOTAs such as MISF and LAMA. While our training utilized the Places2-Standard dataset with 1.8 million images, MISF and LAMA were trained on the Places2-Challenge dataset, which contains 8 million images. Despite using only 22.5% of the images employed by these benchmarks, our model achieved comparable results, showcasing its robustness and efficiency.
Qualitative Comparison. The qualitative comparisons are presented in Fig. 2. For human face samples, M×$\times$×T maintains consistency from the visible regions to the missing regions, such as effectively reconstructing elements like a missing hat. Additionally, M×$\times$×T renders features like eyes with improved fine-grained details, showcasing its strong capability in learning complex representations. In the Places2 dataset, M×$\times$×T effectively captures the spatial layouts in indoor environments and excels at maintaining the architectural integrity of the road surfaces. Such examples highlight our M×$\times$×T has superior spatial perceptions.
Efficiency Comparison. Our M×$\times$×T efficiently reduces spatial dimensions within the transformer and leverages Mamba’s inherent capabilities, both achieving linear complexity. This synergy optimizes operational simplicity while enhancing effectiveness. As shown in Tab. 3, our model comprises 180 million parameters and achieves 110 ms to infer one image, making it suitable for real-time applications.
@@
Table 2: Ablation studies of each component. MB is the Mamba Block with positonal embedding. SRSA is the Spatial Reduced Self-Attention. GDFN is the feed-forward network in [Zamir et al.(2022)Zamir, Arora, Khan, Hayat, Khan, and Yang]. CBFN is the Context Broadcasting Feed-forward Network. Our M×$\times$×T corresponds to configuration (e).
@@ 5.2 Ablation Study
In our comprehensive ablation study conducted on CelebA-HQ, we incrementally enhance the baseline U-Net shape model, observing significant performance improvements with the integration of each component. Results are shown in the Tab. 5.1. The addition of the Mamba Block in configuration (b) and the addition of self-attention in configuration (c) both demonstrated improvement across all evaluated metrics compared to the baseline (a). Notably, self-attention improves in SSIM, suggesting its superior capability in capturing spatial interactions. Mamba showcases the superior in capturing pixel-level interactions, demonstrated by the better PSNR and L1 values. The simultaneous use of Mamba and self-attention in configuration (d) lead to further improvements, indicating that these components effectively complement each other and contribute to a robust model. Configuration (e) is our final model, where we optimize GDFN to CBFN. The overall metrics are further improved. All ablation experiments are trained for 30K iterations. In addition, we followed  [Cui et al.(2022)Cui, Tao, Bing, Ren, Gao, Cao, Huang, and Knoll] to build a light M×$\times$×T with a halved parameter for an efficient evaluation.
5.3 Application: High-Resolution Image Inpainting
Our M×$\times$×T is designed with linear computational complexity, enabling it to effectively handle high-resolution image inpainting task. We directly apply our model, pre-trained on Places2-standard dataset, to real-world high-resolution images, demonstrating its capability. The example is illustrated in Fig. 3.